parking garage fees... The Greenville city Garage charge $5 for the first hour then $2.75 for each additional hour. How long has a car been parked at this garage if the parking charge is $29.75?
Answer by nerdybill(7384)

The Greenville city Garage charge $5 for the first hour then $2.75 for each additional hour. How long has a car been parked at this garage if the parking charge is $29.75?
.
Let h = hours parked
then
5 + 2.75(h-1) = 29.75
2.75(h-1) = 24.75
2.75h-2.75 = 24.75
2.75h = 27.5
h = 10